7d-1=13
can you please tell me all the steps on how to get the correct answer please

Answer by checkley77(12844)

7d-1=13 add 1 to both sides
7d=13+1 add the rioght side
7d=14 divide both sides by 7
d=14/7
d=2 ans.
Proof:
7*2-1=13
14-1=13
13=13